Five times the compliment of an angle less twice the angle's supplement is 40 degrees. Find the measure of the supplement.
Answer by stanbon(48558)

Five times the compliment of an angle less twice the angle's supplement is 40 degrees. Find the measure of the supplement.
------------
Let the angle be "x".
Its complement is "90-x"
Its supplement is "180-x"
----------
EQUATION:
5(90-x)-2(180-x)=40
450-5x-360+2x = 40
-3x = -50
x = 16 2/3 degrees
------------
Supplement = 180-16 2/3 = 163 1/3 degrees
================
Cheers,
Stan H.